# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ H.2.1 Experiments for prediction performance
    • $\alpha_{\text{split}}$ and $\gamma_{\text{split}}$ in (9) are set to 0.95 and 2, respectively, following BART .
@@
    • For $\lambda,$ we reparameterize it to $q_{\lambda},$ where $q_{\lambda}=\pi(\sigma^{2}\leq\hat{\sigma}^{2})$ and $\hat{\sigma}^{2}$ is the variance of the residuals from a linear regression model estimated by the least square method. For the candidates of $q_{\lambda},$ we set $q_{\lambda}=\{0.90,0.95,0.99\}.$ This approach is used in BART .
    • $T_{\max}=\{100,200,300\}$
    • $C_{*}=\{1e-6,1e-4,1e-2,1\}$
    • $\sigma_{\beta}^{2}=\{1e-2,1e-1,1\}$
  • BART (BART )
    • $\alpha_{\text{split}},\gamma_{\text{split}},v,\lambda:$ set to the same values as for ANOVA-BART
@@
    • Further details on the hyperparameters of SSANOVA can be found in gu2014smoothing .
@@
    • Further details on the hyperparameters of MARS can be found in milborrow2017earth .
@@
    • The number of hidden layers =3=3 and the numbers of hidden nodes =(64,32,16)=(64,32,16) (the architecture used in nam ; nbm )
@@ H.2.2 Experiments for component Selection
For the experiment of component selection on synthetic data, the hyperparameters for ANOVA-BART are set to $T_{\max}=300,v=3,q_{\lambda}=0.95$, and $C_{*}=1e-2$.
@@ H.2.3 Experiments for uncertainty quantification
The candidates of the hyperparameters of ANOVA-BART and BART are the same as those used in the prediction performance experiments.

### v3

@@ H.2.1 Experiments for prediction performance
    • $\alpha_{\text{split}}$ and $\gamma_{\text{split}}$ in (6) are set to 0.95 and 2, respectively, following Chipman et al. (2010).
@@
    • For $\lambda,$ we reparameterize it to $q_{\lambda},$ where $q_{\lambda}=\pi(\sigma^{2}\leq\hat{\sigma}^{2})$ and $\hat{\sigma}^{2}$ is the variance of the residuals from a linear regression model estimated by the least square method. For the candidates of $q_{\lambda},$ we set $q_{\lambda}=\{0.90,0.95,0.99\}.$ This approach is used in Chipman et al. (2010).
    • For Madelon data set, we set the burn-in iterations to 2,000 and the sampling iterations to 2,000. For all other data sets, we set the burn-in iterations to 1,000 and the sampling iterations to 1,000.
    • We set the step size for height proposal distribution as $0.01$.
    • $T_{\max}=\{50,100,200,300\}$
    • $C_{*}=\{10^{-6},10^{-4},10^{-2},1\}$
    • $\sigma_{\beta}^{2}=\{10^{-3},10^{-2},10^{-1}\}$
    • $M=\{1,5\}$
  • BART (Chipman et al. (2010))
    • We consider candidate hyperparameters similar to those explored in Chipman et al. (2010).
    • $\alpha_{\text{split}},\gamma_{\text{split}},v,\lambda$, the number of iterations: set to the same values as for ANOVA-BART
@@
    • Further details on the hyperparameters of SSANOVA can be found in Gu (2014).
@@
    • Further details on the hyperparameters of MARS can be found in Milborrow et al. (2017).
@@
    • The number of hidden layers =3=3 and the numbers of hidden nodes =(64,32,16)=(64,32,16) (the architecture used in Agarwal et al. (2021); Radenovic et al. (2022))
@@ H.2.2 Experiments for component Selection
For the experiment of component selection on synthetic data, the hyperparameters for ANOVA-BART are set to $T_{\max}=300,v=3,q_{\lambda}=0.95$, and $C_{*}=10^{-2}$.
@@ H.2.3 Experiments for uncertainty quantification
The candidates of the hyperparameters of ANOVA-BART, BART and MARS are the same as those used in the prediction performance experiments.
H.2.4 Experiments for Stable interpretation
The candidates of the hyperparameters of ANOVA-BART and NAM are the same as those used in the prediction performance experiments.